OpenStudy (hafsahh):

A rocket is fired directly upwards with a velocity of 96 ft/sec, from a platform which is 11 feet above the ground. The equation for the rocket's height, H , as a function of time, t , is given by the function H(t)=-16t^2+96t+11

OpenStudy (zzr0ck3r):

ok

OpenStudy (hafsahh):

I have to find the maximum height

OpenStudy (zzr0ck3r):

Do you know what the vertex of a quadratic?

OpenStudy (zzr0ck3r):

is*

OpenStudy (hafsahh):

the max or min of the function

OpenStudy (zzr0ck3r):

do you know how to find it?

OpenStudy (hafsahh):

not algebraically

OpenStudy (zzr0ck3r):

for the quadratic \(ax^2+bx+c\) the \(x\) coordinate of the vertex is \(x=\dfrac{-b}{2a}\). To find the \(y\) value, which is what you are asked for, plug that into the polynomial.

OpenStudy (hafsahh):

H(t)=-16t^2+96t+11 so H(t)= 16(-b/2a)^2+96(-b/2a)+11?

OpenStudy (hafsahh):

-96/32=-3

OpenStudy (zzr0ck3r):

\(\dfrac{-b}{2a}=\dfrac{-96}{-32}=3\)

OpenStudy (zzr0ck3r):

now plug that into the quadratic to get the y value

OpenStudy (hafsahh):

thanks

OpenStudy (zzr0ck3r):

np